# nonblind {nonblind}

### v1

@@ 3.3 Monotonic Responses and Compliance-Aware Recommendations
Above, we have proposed a potential-outcomes framework for potential agent decisions in response to recommendations, and used this framework to discuss the design of optimal recommendation algorithms in general. We now introduce assumptions that limit the potential outcomes associated with the agent’s decision to simplify the analysis of their endogenous response. We then use these simplifications to shed further light on the structure of optimal recommendations in terms of agent compliance.
In the context of the four response types introduced in the hiring example and repeated in Table 1, manager behavior consistent with the Ignore and Comply responses is easier to explain than behavior consistent with the Defy and Change responses. If the manager believes their private information $U$ is much better than the algorithm’s for some problem instance, they are likely to Ignore the recommendation. Conversely, if they believe the algorithm’s information dominates any additional information in $U$, they will Comply. Using recommendations as a threshold requirement for considering an applicant, as done by hiring managers interviewed by Li et al. (2021), also generates potential outcomes consistent with only Ignore and Comply responses. On the other hand, Defy and Change behaviors are counter-intuitive and typically need bad recommendations, perverse correlation between $X$ and $U$, or adversarial managerial incentives to be justified rationally. Consequently, we now make an assumption to eliminate the potential outcomes associated with Defy and Change responses, similar to the monotonicity assumption of Angrist, Imbens, and Rubin (1996) in the context of instrumental variables and its extensions to multivariate discrete treatments (including Lee and Salanié, 2018).
Assumption 1 (Monotonic response to $f^{\textnormal{rec}}$).
@@
At a high level this assumption restricts the direction of the managers’ response to follow $R$. Hire recommendations ($R=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}$) can turn not-hire decisions into hire decisions ($D=\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}\to D=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}H}}$) but cannot turn hire decisions into not-hire decisions ($D=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}\not\to D=\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}$). Not-hire recommendations ($R=\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}$) do the exact opposite. This restriction exactly rules out the bottom four rows in Table 1. It can be thought of as a restriction on plausible recommendation rules $\mathcal{F}$. Specifically, it rules out that recommendation rules are used that would make sense to Defy (such as a rule that suggests hiring bad candidates) or that would make sense to observe the Change behavior for (such as a rule that only gives a recommendation for bad candidates and never gives one for good candidates).
When the manager’s decision behavior is limited by 1, we can fully describe their behavior for some specific recommendation rule $f^{\textnormal{rec}}$ in terms of their active decisions $D_{\varnothing}=D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}};f^{\textnormal{rec}})$ which they exhibit when not given a recommendation, but the algorithm $f^{\textnormal{rec}}$ is implemented, along with their compliance behavior $C$:
@@
Under 1, the potential outcome triplet $(D_{U}(\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}};f^{\textnormal{rec}}),D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}};f^{\textnormal{rec}}),D_{U}(\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}};f^{\textnormal{rec}}))$ can be fully expressed by the active decision $D_{\varnothing}=D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}};f^{\textnormal{rec}})\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}\}$, which expresses the agent’s decision when not given a recommendation, and the compliance type $C\in\{\textit{Ignore}{},\textit{Comply}{}\}$, which specifies whether the agent adopts $R$ when the recommendation is presented.
Table 2 visualizes 1 by restating the remaining potential-decision types in terms of $(D_{\varnothing},\textit{Comply})$ once 1 is applied. The only information determining the decision $D$ outside of the recommendation $R$ is the active decision $D_{\varnothing}$ and whether the response type $C$ of the manager is Ignore or Comply. The proposition follows directly from an inspection of Table 1. For completeness, a formal proof is given in Appendix A.
$C$ $D_{U}(\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}})$ $D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}})$ $D_{U}({\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}})$
Ignore $D_{\varnothing}$ $D_{\varnothing}$ $D_{\varnothing}$
Comply N $D_{\varnothing}$ H
Table 2: Reduction of potential outcomes under 1. This organization acts as a visual representation of 1, showing how the values that the potential outcomes take reduce to the active decision $D_{\varnothing}\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}{},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}{}\}$ absent a recommendation and the compliance $C\in\{\textit{Ignore}{},\textit{Comply}{}\}$.
@@
$f^{\textnormal{rec}*}=\min_{f:X\mapsto R}$ $\mathbb{E}\left[\ell(Y,R)\>\mathbf{1}\left[C=\textit{Comply}\text{ and }R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]+\ell(Y,D_{\varnothing})\>\mathbf{1}\left[C=\textit{Ignore}\text{ or }R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\right]$ (6)
where $R=f(X)$.
Similarly, this explicit notation allows us to shed further light on the response effect from (5) in terms of compliance and deviations between active decisions $D_{\varnothing}$ and unassisted decisions $D_{0}$:
$\textit{RE}=\mathbb{E}\left[\left(\ell(Y,D_{\varnothing})-\ell(Y,R)\right)\>\mathbf{1}\left[R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}},C=\textit{Ignore}\right]+\left(\ell(Y,D_{\varnothing})-\ell(Y,D_{0})\right)\>\mathbf{1}\left[R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\right]$ (7)
@@
Even with the simplifications introduced by 1, it is not generally possible to learn the optimal recommendation algorithm from past data $(Y,X,D_{0})$ on job performance, characteristics, and unassisted human decisions. This is because neither the compliance behavior $C$ nor the impact of recommendations on active decisions $D_{\varnothing}$ can be learned from such limited data. However, we can now formulate assumptions on these unobserved quantities in order to derive optimal learnable recommendation algorithms in specific cases, thus providing a formal analog and extension of the discussion in Section 3.2:
  • $\circ$
    If compliance is perfect ($C\equiv\textit{Comply}$) and decisions in instances without recommendations are not affected ($D_{\varnothing}\equiv D_{0}$), then the optimal recommendation algorithm is the optimal triage solution $f^{\textnormal{tri}*}$ (which can be learned from data on $(Y,X,D_{0})$).
  • $\circ$
    If compliance is perfect ($C\equiv\textit{Comply}$), and recommendations are limited to be binary ($R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$) or human decisions are dominated by recommended actions (such as when $\mathbb{E}[\ell(Y,D_{\varnothing})|X]\geq\min_{r\in\{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}},\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}\}}\mathbb{E}[\ell(Y,r)|X]$ $X$-a.s. and for all $f^{\textnormal{rec}}\in\mathcal{F}$), then the optimal recommendation algorithm is the optimal decision algorithm $f^{\textnormal{dec}*}$ (which can be learned from data on $(Y,X)$). This applies, specifically, in the case with perfect compliance and no private information of the human decision-maker.
@@
  • $\circ$
    If recommendations do not affect compliance ($C$) or active decisions for instances where no recommendations are given ($D_{\varnothing}=D_{0}$), then the optimal recommendation can generally be seen as the optimal triage decision among compliers only, that is, it solves
    $\min_{f:X\mapsto R}\mathbb{E}\left[\ell(Y,R)\>\mathbf{1}\left[R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]+\ell(Y,D_{0})\>\mathbf{1}\left[R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\middle|C=\textit{Comply}\right]\text{ where }R=f(X).$
@@
  • $\circ$
    Even if the agent takes optimal decisions given their available information, we would require information beyond $(Y,X,D_{0})$, since the optimal recommendation would also depend on the structure and information content of the private signal $U$.
In our experiment below, we combine information about the point distribution of information $(Y,X,U)$ with assumptions on active decisions $D_{\varnothing}$ and compliance $C$ in order to solve for sensible recommendation algorithms.
@@ 4 Complementary Recommendations in an Experiment
We now apply our framework to the design and analysis of recommendation algorithms in a controlled trial. In a pre-registered online experiment, study subjects participate in an incentivized hiring game. [^1] The pre-analysis plan for this experiment can be found at www.socialscienceregistry.org/trials/11857. This empirical application showcases our framework’s ability to map realistic assumptions about agent behavior to the design of practical recommendations that realize complementarity between humans and algorithms.
@@
In the experiment, we collect hypothetical hiring decisions while varying the structure of the recommendation algorithm at the subject level. We apply different assumptions on how subjects will behave within our framework to generate these recommendation algorithms. We find that subjects are able to improve their decisions with algorithms whose recommendations are focused on providing suggestions that are complementary to the subjects’ private information. Moreover, subjects given these focused recommendations perform better than subjects who see naive recommendations based on optimal algorithmic predictions, and outperform decisions by the human or algorithm alone.
4.1 Experimental Design

### v2

@@ 3.3 Monotonic Responses and Compliance-Aware Recommendations
Above, we have proposed a potential-outcomes framework for agent decisions in response to recommendations and used this framework to discuss the design of optimal recommendation algorithms. We now introduce assumptions that limit the potential decision to simplify their analysis. We then use these simplifications to shed further light on the structure of the optimal recommendation algorithm in terms of agent compliance and active decisions.
In the context of the four response types introduced in the hiring example and repeated in Table 1, manager behavior consistent with the Ignore and Comply responses is easier to make sense of than behavior consistent with the Defy and Change responses. If the manager believes the algorithm made a bad recommendation given their private information $U$ about the applicant, they are likely to Ignore the recommendation. Conversely, if they believe the algorithm makes a good recommendation given their private information $U$ about the applicant, they will Comply. Using recommendations as a threshold requirement for considering an applicant, as done by hiring managers interviewed by Li et al. (2021), also generates potential outcomes consistent with only Ignore and Comply responses. On the other hand, Defy and Change behaviors are counter-intuitive and typically need bad recommendation algorithm designs, perverse correlation between $X$ and $U$, or adversarial managerial incentives to be justified rationally. Consequently, we now make an assumption to eliminate the potential outcomes associated with Defy and Change responses, similar to the monotonicity assumption of Angrist, Imbens, and Rubin (1996) in the context of instrumental variables and its extensions to multivariate discrete treatments (including Lee and Salanié, 2018).
Assumption 1 (Monotonic response to $R$).
@@
At a high level, this assumption restricts the direction of the managers’ response to follow the algorithm’s recommendation for the applicant. Hire recommendations ($R=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}$) can turn not-hire decisions into hire decisions ($D=\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}\to D=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}H}}$) but cannot turn hire decisions into not-hire decisions ($D=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}\not\to D=\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}$). Not-hire recommendations ($R=\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}$) do the exact opposite.
The monotonicity restriction exactly rules out the bottom four rows of Table 1. It can be considered a restriction on plausible recommendation rules $\mathcal{F}$. Specifically, it rules out that recommendation rules are used that would make sense to Defy (such as a rule that suggests hiring bad candidates) or that would make sense to observe the Change behavior for (such as a rule that only gives a recommendation for bad candidates and never gives one for good candidates).
When the manager’s decision behavior is limited by 1, we can fully describe their behavior for some specific recommendation rule $f^{\textnormal{rec}}$ in terms of their active decisions $D^{\varnothing}_{U}=D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}};f^{\textnormal{rec}})$ which they exhibit in response to the algorithm $f^{\textnormal{rec}}$ being implemented, along with their compliance behavior $C_{U}$:
@@
Under 1, the potential outcome triplet $(D_{U}(\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}};f^{\textnormal{rec}}),D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}};f^{\textnormal{rec}}),D_{U}(\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}};f^{\textnormal{rec}}))$ can be fully expressed by the active decision $D^{\varnothing}_{U}(f^{\textnormal{rec}})=D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}};f^{\textnormal{rec}})\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}\}$, which expresses the agent’s decision when not given a recommendation, and the compliance type $C_{U}(f^{\textnormal{rec}})\in\{\textit{Ignore}{},\textit{Comply}{}\}$, which specifies whether the agent adopts $R=f^{\textnormal{rec}}(X)$ when the recommendation is presented.
Table 2 visualizes 1 by restating the remaining potential-decision types in terms of $(D^{\varnothing}_{U},C_{U})$ once 1 is applied. The only information determining the decision $D$ outside of the recommendation $r$ is the active decision $D^{\varnothing}_{U}$ and whether the response type $C_{U}$ of the manager is Ignore or Comply. The proposition follows directly from an inspection of Table 1. For completeness, a formal proof is given in Appendix A.
$C_{U}$ $D_{U}(\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}})$ $D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}})$ $D_{U}({\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}})$
Ignore $D^{\varnothing}_{U}$ $D^{\varnothing}_{U}$ $D^{\varnothing}_{U}$
Comply N $D^{\varnothing}_{U}$ H
Table 2: Reduction of potential outcomes under 1. This organization acts as a visual representation of 1, showing how the values that the potential outcomes take reduce to the active decision $D^{\varnothing}_{U}\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}{},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}{}\}$ absent a recommendation and the compliance $C_{U}\in\{\textit{Ignore}{},\textit{Comply}{}\}$.
@@
$f^{\textnormal{rec}*}=\underset{f:X\mapsto R}{\mathrm{argmin}}$ $\mathbb{E}\left[\ell(Y,R)\>\mathbf{1}\left[C_{U}=\textit{Comply}\text{ and }R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]+\ell(Y,D^{\varnothing}_{U})\>\mathbf{1}\left[C_{U}=\textit{Ignore}\text{ or }R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\right]$ (6)
where $R=f(X)$
Similarly, this explicit notation allows us to shed further light on the response effect from (5) in terms of compliance and deviations between active decisions and unassisted decisions:
$\textit{RE}=\mathbb{E}\left[\left(\ell(Y,D^{\varnothing}_{U})-\ell(Y,R)\right)\>\mathbf{1}\left[R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}},C_{U}=\textit{Ignore}\right]+\left(\ell(Y,D^{\varnothing}_{U})-\ell(Y,D^{0}_{U})\right)\>\mathbf{1}\left[R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\right]$
@@
Even with the simplifications introduced by 1, it is not generally possible to learn the optimal recommendation algorithm from past data $(Y,X,D^{0}_{U})$ on outcomes, characteristics, and unassisted human decisions. This is because the joint distributions of active decisions and compliance with outcomes and characteristics, $(Y,X,D^{\varnothing}_{U},C_{U})$ cannot be identified from such limited data without further assumptions. However, we can now formulate conditions on these unobserved quantities to derive optimal learnable recommendation algorithms in specific cases, thus providing a formal analog and extension of the discussion in Section 3.2:
  • If compliance is perfect ($C_{U}\equiv\textit{Comply}$) and decisions in instances without recommendations are not affected ($D^{\varnothing}_{U}\equiv D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}},f^{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}})$), then the optimal recommendation algorithm is the optimal triage solution $f^{\textnormal{tri}*}$ (which can be learned from data on $(Y,X,D^{0}_{U})$).
  • If compliance is perfect ($C_{U}\equiv\textit{Comply}$), and recommendations are limited to be binary ($R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$) or human decisions are dominated by recommended actions (such as when $\mathbb{E}[\ell(Y,D^{\varnothing}_{U})|X]\geq\min_{r\in\{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}},\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}\}}\mathbb{E}[\ell(Y,r)|X]$ for all $f^{\textnormal{rec}}\in\mathcal{F}$), then the optimal recommendation algorithm is the optimal decision algorithm $f^{\textnormal{dec}*}$ (which can be learned from data on $(Y,X)$). This applies, specifically, in the case with perfect compliance and no private information of the human decision-maker.
@@
  • If the recommendation algorithm is independent of compliance and active decisions ($D^{\varnothing}_{U}\equiv D^{0}_{U}$), then the optimal recommendation can generally be seen as the optimal triage decision among compliers only,
    $\underset{f:X\mapsto R}{\mathrm{argmin}}\ \mathbb{E}\left[\ell(Y,R)\>\mathbf{1}\left[R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]+\ell(Y,D^{0}_{U})\>\mathbf{1}\left[R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\middle|C_{U}=\textit{Comply}\right]\text{ where }R=f(X).$
@@
  • Even if the agent makes optimal decisions given their available information, we would require information beyond $(Y,X,D^{0}_{U})$ since the optimal recommendation would also depend on the structure and information content of the private signal $U$.
In our experiment below, we combine information about the point distribution of information $(Y,X,U)$ with assumptions on active decisions $D^{\varnothing}_{U}$ and compliance $C_{U}$ in order to solve for sensible recommendation algorithms.
@@ 4 Complementary Recommendations in an Experiment
We now apply our framework to the design and analysis of recommendation algorithms in a controlled trial. In a pre-registered online experiment, study subjects participate in an incentivized hiring game. [^2] The pre-analysis plan for this experiment can be found at www.socialscienceregistry.org/trials/11857. This empirical application showcases our framework’s ability to map realistic assumptions about agent behavior to the design of practical recommendation algorithms that realize complementarity between humans and algorithms.
@@
In the experiment, we collect hypothetical hiring decisions while varying the structure of the recommendation algorithm at the subject level. We apply different assumptions on how subjects will behave within our framework to generate these recommendation algorithms. We find that subjects, on average, improve their decisions with algorithms whose recommendations are focused on providing suggestions that are complementary to the subjects’ private information. Moreover, subjects given these focused recommendations perform better than subjects who see naive recommendations based on optimal algorithmic predictions and outperform decisions by the human or algorithm alone.
4.1 Experimental Setup